When reading a research article, you encounter the following statement: “The t-test indicated a significant difference...

  1. When reading a research article, you encounter the following statement: “The t-test indicated a significant difference between the mean for men (M = 5.4) and for women (M = 9.3), with df(58) = 7.93, p ˂ .05.
    1. What is the N in this study?_______________
    2. What statistical procedure was used?______________________
    3. What is the probability of a Type I error?____________________
Answer #1

Answer

(A) we know that relationship between degree of freedom and sample size N is given as

N = df + 2

we have df = 58

So, required value of N = 58 +2 = 60

Therefore, N in this study is 60

(B) It is clear that we have men and women in this hypothesis testing with t test, so it will be an independent samples t test where two samples are men and women. Inferential statistical procedure used is independent samples t test for mean difference testing.

(C) it is given in the statement that p < 0.05, this means we are using 0.05 level of significance in this hypothesis testing. We know that the probability of type I error is always equal to the significance level, so we can say that the probability of type I error is 0.05.